The hypotenuse of a right triangle is (c) and one (a)leg is 24 and other (b)leg is 32.
Find the sin A and tan A
also
a= 2.8m, B= 41.8 degrees
and
The hypotenuse of a right triangle is 46.0in and one leg is 35.4 Find the length of the other leg.

Answer by RAY100(1637)

right triangle means pythagorous theorem holds,,,,,c^2=a^2+b^2
finding c from given two legs,
,
c^2 = 24^2 + 32^2 =1600
c=40
,
rough sketch of triangle helps. Label vertex opposite side with same letter, but capital.
,
sin A = opposite/hyp = 24 /40 = .6,,,,,arc sin a= 36.9 deg
,
tan a = opposite / adjacent = 24/32 = .75, arc tan a = 36.9 deg
,
other problem
again use c^2 = a^2 + b^2
46^2 = 35.4^2 + b^2
b^2 = 46^2-35.4^2
b^2=862.8
b = 29.4
,
,Last problem
again, a sketch helps keep everything straight
sum of angles = 180, therefore find last angle, 180-90-41.8 =48.2 deg
sin of 48.2 = 2.8 / c
c=2.8 / sin 48.2 = 3.756
,
sin 41.8 = b/3.756
b=3.756 * sin 41.8 = 2.5
therefore all angles and all sides are labeled
,
,
,
,